# A quantum implementation of high-order power method for estimating geometric entanglement of pure states

### v1

@@ 2.3 Simulation Results
Since GHZ[$9$] is calibrated against itself, to demonstrate our mitigation technique on realistic noise models we show only the state $Random[6]$ ($p$ estimated with GHZ[$6$]). The graph for GHZ[$9$] is in Supplemental Figure 6(f). In Fig. 5(a) we contrast the classical HOPM to FakeLima’s $\overline{E}_{G}\approx 0.685$ with $\sigma_{E_{G}}\approx 0.034$ and FakeSherbrooke’s $\overline{E}_{G}\approx 0.682$ with $\sigma_{E_{G}}\approx 0.028$. For FakeLima we estimated the error rate to be $p\approx 24.68\times 10^{-4}$ and after mitigation we see in Fig. 5(b) that $E_{G}^{\prime}\approx 0.641$; for FakeSherbrooke we estimate the error rate to be $p\approx 25.18\times 10^{-4}$ and compute the mitigated value $E_{G}^{\prime}\approx 0.636$. See Supplementary Materials for more states simulations and mitigation results for these noise models.
Figure 5: Simulation results of QHOPM for Random[$6$] with more realistic noise models and its mitigation. (a) and (b) show unmitigated and mitigated results for Random[$6$], respectively. Each simulation was run 10 times (with the same 10 random initial separable states) with $1\times 10^{5}$ shots per each measurement. Solid lines represent the mean geometric entanglement, $E_{G}{}$, and the shaded colours represent the standard deviation. The colour of the lines represents the noise model used for the simulation (see legend). Red dotted lines show results of classical HOPM.
@@ 3 Discussion
In this work we presented QHOPM, an iterative quantum algorithm for the approximation of the geometric measure of entanglement of pure states on near term quantum devices as well as a noise mitigation technique to improve its experimental resolution. QHOPM is the quantum adaptation of the well-known classical HOPM approach [20, 21, 22, 26, 27, 29] for rank-1 tensor approximation.
@@
While QHOPM does not converge any differently than classical HOPM, its potential utility lies in the observation that it requires only $n+1$ qubits to compute the GE of a quantum state (on $n$ qubits) while an exponential data structure (and so exponential runtime) is assumed by classical HOPM (for quantum states). We also show that the number of measurements required, per algorithm iteration, scales linearly with $n$.
@@
We showed (see Remarks 8 and 9) that the effect of depolarizing noise on QHOPM for the studied noise rates ($p\leq 0.05$) does not affect its convergence, but rather shifts the final result. The same observation was made in simulations using more realistic noise models implemented in IBM Qiskit’s FakeLima and FakeSherbrooke simulation backends. This allowed us to assume and successfully apply the mitigation technique developed for the DN model on these unknown noise channels. The noise rate for the unknown models was estimated using GHZ as a reference state with a known GE.
@@
Another mitigation approach is presented in [46], where the authors assumed that the channel is applied only on the final stage before each measurement with the “total” error rate $p_{tot}$, which they propose to recover from measuring $\operatorname{Tr}(\rho^{2})$. This approach is equivalent to ours if we use $p_{tot}=1-q^{d}$ and treat $p$ and $d$ as parameters (see Remark 9).
@@
Noise analysis was not the main topic of this work and the assumptions we made are simplified. A more consistent and rigorous approach would be to consider application of DN channel to individual gates with different noise rates or to study the noise model of a particular quantum device.
@@
To our knowledge, the only other quantum algorithms for calculating geometric entanglement are variational algorithms where the ansatz is a separable state [32, 31]. We compare one of these, the VDGE algorithm [31], with QHOPM. VDGE converges much more slowly than QHOPM, needing on the order of $10^{2}$ iterations on GHZ compared to QHOPM’s $10^{0}$. The authors of VDGE claim that the optimizer they use (complex simultaneous perturbation stochastic approximation) is robust against noise. It gives results within $\epsilon\approx 10^{-4}$ absolute accuracy for the three qubit random state case, the error gradually grows as the number of qubits increases (e.g.$\epsilon\approx 10^{-2}$ for six qubit random states). We found that on a random 6 qubit state with realistic noise models and a relatively simple mitigation scheme, QHOPM got results within the same range $\epsilon\approx 10^{-2}$. Like QHOPM, these variational approaches are sensitive to the initial separable state and one is never sure if the result obtained is a local minima or the global minimum. Also, like any quantum variational approach, VDGE is prone to have barren plateaus as the number of qubits grow, which makes it hard to implement on large scale devices.
@@ 3.1 Quantum circuit simulation
To investigate the effect of noise on our algorithm, we use the depolarizing noise (DN) model [49] and more realistic noise models implemented in “FakeLima” and “FakeSherbrooke” backends, developed to mimic real IBM hardware. We let the parameter of the DN model (referred to in this work as “noise rate” or “error rate”) $p\in\{0,0.001,0.01,0.05\}$ and apply DN to all one-, two- and three-qubit gates used in our circuits via the Qiskit AerSimulator.
@@
  • 3, 6, and 9 qubit GHZ states, each with known $E_{G}=0.5$;
@@
  • 3, 6 and 9 qubit random states.
@@
3.2 Depolarizing noise mitigation
@@
where $I$ is a $2$-dimensional identity matrix; $0\leq p\leq 4^{n}/(4^{n}-1)$ is a parameter of the model (which we will refer to as noise rate). When $p=1$ the channel is fully depolarizing. When $p=4^{n}/(4^{n}-1)$ the channel has with (uniformly) random Pauli errors on each qubit.
Typically the DN channel is applied to each gate in a quantum circuit individually and with different rates. This leads to cumbersome analysis and a strong dependence on the structure of the circuit under study. To avoid this here we consider the following simplification: we assume that the DN channel is applied $d$ times at a constant rate $p$ to all qubits after each layer of gates, where $d$ is the circuit depth (maximum number of operations on a qubit in the circuit). The DN channel commutes with any unitary operation:
@@
Having the DN channel defined by Eq. (15) in terms of the Hadamard test procedure, the measured $P_{i}(s)$ and $\lambda^{(k)}$ is expressed as
$P_{i}(s)\approx\left(1+\eta P_{i}^{\prime}(s^{\perp})+\ldots\right)\,P_{i}^{\prime}(s);$ (16)
@@
where $\eta=(1-q^{2})$; $P_{i}^{\prime}(s)$, $P_{i}^{\prime}(s^{\perp})$ and $\lambda^{\prime(k)}$ are the noise-free values of $P_{i}(s)$, $P_{i}(s^{\perp})$ and $\lambda^{(k)}$, respectively; dots in Eq. (16) correspond to the terms of the second order and higher in $\eta$; $\gamma^{\prime}$ is the phase of the noise-free scalar product $\bra{\bf 0}{V^{(k)}}^{\dagger}U_{\psi}\ket{\bf 0}=\lambda^{\prime(k)}e^{-i\gamma^{\prime}}$, which is recovered from
$\tan\gamma^{\prime}=q\frac{\langle Y\rangle_{a}}{\langle X\rangle_{a}}.$ (18)
@@
To obtain $P_{i}(s)$, defined by Eq. (10), we need to measure $u_{i,s}^{(k)}$, which in terms of the Hadamard test procedure is done by measuring
$\begin{split}&\langle X\rangle_{a}=(P^{(a)}(-)-P^{(a)}(+))=1-2P^{(a)}(+),\\ &\langle Y\rangle_{a}=1-2P^{(a)}(-i),\end{split}$ (19)
@@
are the “noisy” probabilities of the ancilla qubit being in the state $\ket{+}$ and |$\ket{-i}$, respectively, and ${P^{(a)}}^{\prime}(+)$ and ${P^{(a)}}^{\prime}(-i)$ are their “non-noisy” counterparts. After substituting Eqs. (20) to Eqs. (19), for $u_{i,s}^{(k)}$ we have:
$|u_{i,s}^{(k)}|^{2}=\langle X\rangle_{a}^{2}+\langle Y\rangle_{a}^{2}=q^{2d}[1-(1-q^{2})\sin^{2}\gamma^{\prime}]|u_{i,s}^{\prime(k)}|^{2},$ (21)
where $u_{i,s}^{\prime(k)}$ is a non-noisy value of $u_{i,s}^{(k)}$; $\gamma^{\prime}$ is the phase of $u_{i,s}^{\prime(k)}=e^{i\gamma^{\prime}}|u_{i,s}^{\prime(k)}|$, which originates from the global phase and local $z$-rotations. We know (see Remark 3) that in the absence of noise, the global phase does not affect the convergence of the algorithm, however since $d$ for $\langle Y\rangle_{a}^{\prime}=(\lambda^{\prime(k)}\sin\gamma^{\prime})$ is larger, in the noisy case we will have additional terms in the sum in Eq. (21). In the experiment $\gamma^{\prime}$ is recovered from
$\tan\gamma^{\prime}=\frac{\langle Y\rangle_{a}^{\prime}}{\langle X\rangle_{a}^{\prime}}=q\frac{\langle Y\rangle_{a}}{\langle X\rangle_{a}}.$
@@
$|u_{i,s^{\perp}}^{(k)}|^{2}=q^{2d+2}[1-(1-q^{2})\sin^{2}\gamma^{\prime}]|u_{i,s^{\perp}}^{\prime(k)}|^{2},$ (22)
where the angle $\gamma^{\prime}$ is the same as for the state |$\ket{s}$, since $U_{s^{\perp}}=U_{s}X$ does not change the phase [^3] For real hardware or more realistic simulations, the phase might change depending upon the implementation of $X$, however, it will not change $P_{i}(s)$ drastically..
Substituting Eqs. (21) and (22) to Eq. (10), we have for $P_{i}(s)$:
$P_{i}(s)=\frac{|u^{\prime}_{i,s}|^{2}}{|u^{\prime}_{i,s}|^{2}+|u^{\prime}_{i,s^{\perp}}|^{2}-\eta|u^{\prime}_{i,s^{\perp}}|^{2}}\approx\left(1+\eta P_{i}(s^{\perp})+\ldots\right)P_{i}(s),$ (23)
where $\eta=(1-q^{2})$; the right-hand side is the Taylor expansion of the left side near $\eta=0$; $P_{i}(s^{\perp})$ is a non-noisy value of $P_{i}(s^{\perp})$. This proves the first statement of the Proposition.
@@
$\begin{split}\lambda^{{(k)}^{2}}&=\langle X\rangle_{a}^{2}+\langle Y\rangle_{a}^{2}=q^{2d}[1-\eta\sin^{2}\gamma^{\prime}]\lambda^{\prime{(k)}^{2}},\end{split}$ (24)
@@
Remark 8.
In terms of the Hadamard test procedure, assuming a fixed initial separable state, small noise rates $p\ll 1$ do not affect convergence of QHOPM and contribute mainly to the shift of entanglement eigenvalue, defined by Eq. (17).
Indeed, sufficiently small $p$ will result in $\eta\ll 1$, which allows to neglect the terms with $\eta$ of the first order and higher in Eq. (16) having the main contribution to be the shift in $\lambda$, defined by Eq. (17). The convergence of QHOPM is defined by line 2 in Algorithm 2, where $\lambda^{\prime(k)}$ and $\lambda^{\prime(k-1)}$ change by coefficients different by at most $q=1-p\approx 1$. Putting the accuracy $\epsilon^{\prime}=(q^{d}\epsilon)$ or having a fixed sufficiently large number of iterations allows to avoid the effect of noise on convergence of QHOPM, leaving only the effect of shift in $\lambda$. Note, that this assumes a fixed initial separable state; different initial states might have different convergence.
In this work, the largest error rate we have studied is $p=0.05$ which gives $\eta\approx 0.1$. Thus, according to Proposition 7 and assuming $P_{i}^{\prime}(s)=P_{i}^{\prime}(s^{\perp})=0.5$, $P_{i}(s)$ changes by at most a factor of approximately $1.025$. We neglect this effect in what follows and take into account only the changes of $\lambda^{{(k)}}$, which even for relatively shallow $U_{\psi}$ may have a significant effect. With these assumptions we approximate the true $E_{G}^{\prime}$ as follows:
@@
Proposition 7, Remark 8 and Eq. (25) are valid for a DN channel with the noise rate $p_{tot}=1-q^{d}$ applied once in the end of the circuit just before the measurements.

### v2

@@ 2.3 Simulation Results
Since GHZ[$9$] is calibrated against itself, to demonstrate our mitigation technique on realistic noise models we show only the state Random[$6$] ($p$ estimated with GHZ[$6$]). The graph for GHZ[$9$] is in Supplemental Figure 7(f). In Fig. 6 we show the median absolute error of the QHOPM results for Random[$6$] state within (a) FakeSherbrooke and (c) FakeLima backends, and their mitigation results (b) and (d), respectively. See Supplementary Materials for more state simulations and mitigation results for these noise models.
Figure 6: The error ($\text{m}(\mathcal{F}_{k})-\hat{E}_{G}$) of QHOPM approximating $E_{G}$ for 100 different samples of Random[$n$] with realistic noise models and their mitigation. (a) and (b) show unmitigated and mitigated results, respectively, for Random[$n$] with $n\in\{3,4,5,6\}$, simulated with the FakeSherbrooke Qiskit backend. (c) and (d) show the same for the FakeLima Qiskit backend ($n\in\{3,4\}$ since FakeLima imitates a 5 qubit device). Solid lines represent the median error over the initial separable states and the shaded colours represent the interquartile range. The colour of the lines represents the number of qubits (see legend).
@@ 3 Discussion
In this work we presented QHOPM, an iterative quantum algorithm for the approximation of the geometric measure of entanglement of pure states on near term quantum devices as well as a simple noise mitigation technique to improve its experimental resolution. QHOPM is the quantum adaptation of the well-known classical HOPM approach [20, 21, 22, 26, 27, 29] for rank-1 tensor approximation.
@@
While QHOPM does not converge differently than classical HOPM and gives similar results (assuming a sufficient amount of shots and relatively low noise rate) its potential utility lies in the observation that it requires only $n+1$ qubits to compute the GE of a quantum state (on $n$ qubits) while classical HOPM naively requires an exponential data structure (and so exponential runtime) to work with quantum states. We also show that the number of measurements required per algorithm iteration scales linearly with $n$. In our work, to avoid the problem of post-selection we have assumed Hadamard test-based measurements, which have recently been successfully implemented on real devices [46].
@@
We observed that the effect of depolarising noise on QHOPM for the studied noise rates ($p\leq 0.05$) and number of shots ($10^{5}$) does not affect its convergence, but rather shifts the final result. The same observation was made in simulations using more realistic noise models implemented in IBM Qiskit’s FakeLima and FakeSherbrooke simulation backends. This allowed us to assume and successfully apply the mitigation technique developed for the DN model on these unknown noise channels. The noise rate for the unknown models was estimated using GHZ[$n$] as a reference state with a known GE.
@@
Another mitigation approach is presented in [47], where the authors assumed that the DN channel is applied only on the final stage before each measurement with the “total” error rate $p_{tot}$, which they propose to recover from measuring $\operatorname{Tr}(\rho^{2})$. This approach is equivalent to ours if we use $p_{tot}=1-q^{d}$ and treat $p$ and $d$ as parameters (see Remark 8).
@@
We note that noise analysis was not the main topic of this work and the assumptions we made are simplified. A more consistent and rigorous approach would be to consider the application of the DN channel with different noise rates to individual gates twirled over the Haar distribution (using, say, Clifford 2-design) [48, 49] or to study the noise model of a particular quantum device.
@@
To our knowledge, the only other quantum algorithms for calculating geometric entanglement are variational algorithms where the ansatz is a separable state [32, 31]. We now compare one of these, the VDGE algorithm [31], with QHOPM. VDGE converges more slowly than QHOPM, needing on the order of $10^{2}$ iterations compared to QHOPM’s $10^{0}$, with noise or without. In a noiseless simulation (for Haar random target states) with 3 qubits, the median absolute error of VDGE’s best results for each target state converges to approximately $10^{-4}$ (see Fig. 3 in [31]). As the number of qubits increases to six, the error grows to approximately $10^{-2}$. In similar simulations with noiseless QHOPM, the accuracy of the results seems less dependent on the number of qubits and more on the number of shots, satisfying the Chernoff bound (see Fig. 4). So in the noiseless case, while VDGE is more accurate with fewer shots per measurement on small numbers of qubits, QHOPM can improve its accuracy by increasing the number of shots, independent of the number of qubits (at least, for the range of qubits studied). When realistic noise is introduced, the accuracy of QHOPM’s results (see Fig. 6) are of the order of VDGE results (see Fig. 6 [31]). The authors of VDGE, use an optimizer (complex simultaneous perturbation stochastic approximation) that is claimed to be robust against noise. We note that, in the presence of noise, QHOPM still converges to a value and simple mitigation techniques recover a number closer to the noiseless value. Like QHOPM, these variational approaches are sensitive to the initial separable state and one is never sure if the result obtained is a local minimum or the global minimum. Also, like any quantum variational approach, VDGE is prone to have barren plateaus as the number of qubits grows, which makes it difficult for the algorithm to find a good solution. In all examples we have studied, QHOPM converges in approximately 4 iterations, however we have no formal statement about the convergence rate of the algorithm. The convergence rate of QHOPM will be a complex extension of the known HOPM convergence rate for real valued tensors [43, 44].
@@ 3.1 QHOPM simulation
To investigate the effect of noise on our algorithm, we use the depolarising noise (DN) model [52] and more realistic noise models implemented in the “FakeLima” and “FakeSherbrooke” Qiskit backends, developed to mimic real IBM hardware. To simplify the analysis, we let the parameter of the DN model (referred to in this work as “noise rate” or “error rate”) $p\in\{0,0.001,0.01,0.05\}$ and apply DN to all one-, two- and three-qubit gates used in our circuits via the Qiskit AerSimulator.
@@
  • 3, 4, 5, 6, and 9 qubit GHZ states, each with known $E_{G}=0.5$;
@@
  • 3, 4, 5, 6 and 9 qubit random states.
@@
3.2 Depolarising noise mitigation
@@
where $I$ is a $2$-dimensional identity matrix; $0\leq p\leq 4^{n}/(4^{n}-1)$ is a parameter of the model (which we will refer to as noise rate). When $p=1$ the channel is fully depolarising. When $p=4^{n}/(4^{n}-1)$ the channel is equivalent to random Pauli errors applied (uniformly) on each qubit with equal probability.
A more consistent approach would be to “convert” the noise channel in the system to depolarising noise using, say, Clifford 2-design [48, 49] and assume different noise rates for different types of gates. As our work was not aimed at detailed noise analysis we consider the following simplification: we assume that the DN channel is applied $d$ times at a constant rate $p$ to all qubits after each layer of gates, where $d$ is the circuit depth (maximum number of operations on a qubit in the circuit). The DN channel commutes with any unitary operation:
@@
Given the DN channel defined by Eq.  (16), in terms of the Hadamard test procedure and sufficient number of shots, the measured $P_{i}(s)$ ($P_{i}(s,U)$ for some arbitrary $U$) and $\lambda^{(k)}$ can be expressed as
$P_{i}(s)=\frac{P_{i}^{\prime}(s)}{1-\eta P_{i}^{\prime}(s^{\perp})}\approx\left(1+\eta P_{i}^{\prime}(s^{\perp})+\ldots\right)\,P_{i}^{\prime}(s);$ (17)
@@
where $\eta=(1-q^{2})>0$; $P_{i}^{\prime}(s)$, $P_{i}^{\prime}(s^{\perp})$ and $\lambda^{\prime(k)}$ are the noise-free values of $P_{i}(s)$, $P_{i}(s^{\perp})$ and $\lambda^{(k)}$, respectively; dots in Eq.  (17) correspond to the terms of the second order and higher in $\eta$; $\gamma^{\prime}$ is the phase of the noise-free scalar product $\bra{\bf 0}{V^{(k)}}^{\dagger}U_{\psi}\ket{\bf 0}=\lambda^{\prime(k)}e^{-i\gamma^{\prime}}$, which is recovered from
$\tan\gamma^{\prime}=q\frac{\langle Y_{a}\rangle}{\langle X_{a}\rangle}.$ (19)
@@
To obtain $P_{i}(s)$, defined by Eq.  (11), we need to measure $|\braket{b_{[i]}}{w}|$ |, which in terms of the Hadamard test procedure is done by measuring
$\begin{split}&\langle X_{a}\rangle=1-2P^{(a)}(+),\\ &\langle Y_{a}\rangle=1-2P^{(a)}(-i),\end{split}$ (20)
@@
are the “noisy” probabilities of the ancilla qubit being in the state $\ket{+}$ and |$\ket{-i}$, respectively, and ${P^{(a)}}^{\prime}(+)$ and ${P^{(a)}}^{\prime}(-i)$ are their “non-noisy” counterparts. After substituting Eq.  (21) to Eq.  (20), for $u_{i,s}^{(k)}$ we have:
$|\braket{b_{[i]}}{w}|^{2}=\langle X_{a}\rangle^{2}+\langle Y_{a}\rangle^{2}=q^{2d}[1-(1-q^{2})\sin^{2}\gamma^{\prime}]|\braket{b_{[i]}}{w}^{\prime}|^{2},$ (22)
where $\braket{b_{[i]}}{w}^{\prime}$ is a non-noisy value of $\braket{b_{[i]}}{w}$; $\gamma^{\prime}$ is the phase of $\braket{b_{[i]}}{w}^{\prime}=e^{i\gamma^{\prime}}|\braket{b_{[i]}}{w}^{\prime}|$ |, which originates from the global phase and local $z$-rotations. We know (see Remark 3) that in the absence of noise, the global phase does not affect the convergence of the algorithm, however since $d$ for $\langle Y_{a}\rangle^{\prime}=(\sin\gamma^{\prime}|\braket{b_{[i]}}{w}^{\prime}|)$ is larger, in the noisy case we will have additional terms in the sum in Eq.  (22). In the experiment $\gamma^{\prime}$ is recovered from
$\tan\gamma^{\prime}=\frac{\langle Y_{a}\rangle^{\prime}}{\langle X_{a}\rangle^{\prime}}=q\frac{\langle Y_{a}\rangle}{\langle X_{a}\rangle}.$
@@
$|\braket{b^{\perp}_{[i]}}{w}|^{2}=q^{2d+2}[1-(1-q^{2})\sin^{2}\gamma^{\prime}]|\braket{b^{\perp}_{[i]}}{w}^{\prime}|^{2},$ (23)
where the angle $\gamma^{\prime}$ is the same as for the state |$\ket{b}$, since $U_{b^{\perp}}=U_{b}X$ does not change the phase [^3] For real hardware or more realistic simulations, the phase might change depending upon the implementation of $X$, however, it will not change $P_{i}(s)$ drastically..
Substituting Eq.  (22) and Eq.  (23) to Eq.  (11), we have for $P_{i}(s)$:
$\begin{split}P_{i}(s)&=\frac{|\braket{b_{[i]}}{w}^{\prime}|^{2}}{|\braket{b_{[i]}}{w}^{\prime}|^{2}+|\braket{b^{\perp}_{[i]}}{w}^{\prime}|^{2}-\eta|\braket{b^{\perp}_{[i]}}{w}^{\prime}|^{2}}\\ &=\frac{P^{\prime}_{i}(s)}{1-\eta P^{\prime}_{i}(s^{\perp})}\approx\left(1+\eta P^{\prime}_{i}(s^{\perp})+\ldots\right)P^{\prime}_{i}(s),\end{split}$ (24)
where $\eta=(1-q^{2})>0$; the right-hand side is the Taylor expansion of the left side near $\eta=0$; $P^{\prime}_{i}(s^{\perp})$ is a non-noisy value of $P_{i}(s^{\perp})$. This proves the first statement of the Proposition.
@@
$\begin{split}\lambda^{{(k)}^{2}}&=\langle X_{a}\rangle^{2}+\langle Y_{a}\rangle^{2}=q^{2d}[1-\eta\sin^{2}\gamma^{\prime}]\lambda^{\prime{(k)}^{2}},\end{split}$ (25)
@@
If all the $\eta$ terms in Eq.  (17) are negligible ($p\ll 1$), the one-qubit tomography procedure $\mathcal{T}_{i}$ will return angles with errors of similar magnitude for any circuit depth. If these errors can be neglected, the convergence condition of QHOPM, given in line 2 in Algorithm 2, will change only by the prefactor defined in Eq.  (18). In this work, the largest error rate that we have studied is $p=0.05$ which gives $\eta\approx 0.1$. Thus, according to Proposition 7 and assuming $P_{i}^{\prime}(s)=P_{i}^{\prime}(s^{\perp})=0.5$, $P_{i}(s)$ changes at most by approximately $0.05$. This change does not affect the convergence, as it is seen from the simulations. We neglect this effect in what follows and take into account only the changes of $\lambda^{{(k)}}$, which even for relatively shallow $U_{\psi}$ may be significant. With these assumptions we approximate the true $E_{G}^{\prime}$ as follows:
@@
Proposition 7 and Eq.  (26) are valid for a DN channel with the noise rate $p_{tot}=1-q^{d}$ applied once in the end of the circuit just before the measurements.